# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v1

@@ Appendix C Omitted Proofs
$W_{\mathcal{G}^{\prime}}(\bm{\pi})=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi_{j}}w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}+\gamma_{j,\pi_{j}}\right)\geq\sum_{j=1}^{N}\frac{(\mu_{\pi_{j}}-\Delta)w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}=\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}-k\Delta\geq\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}-K\Delta.$ (56)
@@
$W_{\mathcal{G}^{\prime}}(\tilde{\bm{\pi}})-W_{\mathcal{G}^{\prime}}(\bm{\pi})\geq\sum_{k^{\prime}=1}^{\tilde{k}}\mu_{k^{\prime}}-K\Delta-\left(\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}+K\Delta+N\Gamma\right)\geq\mu_{\tilde{k}}-2K\Delta-N\Gamma.$ (57)
@@
$W_{\mathcal{G}^{\prime}}(\tilde{\bm{\pi}})-W_{\mathcal{G}^{\prime}}(\bm{\pi})\geq\mu_{\tilde{k}}-2K\Delta-N\Gamma\geq\delta-2K\cdot\frac{\delta}{4K}-N\cdot\frac{\delta}{4N}=\frac{\delta}{4}>0.$ (58)
@@
$\,\mathbb{P}\left[\exists\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N},\bm{\pi}^{\prime}\text{ and }\bm{\pi}^{\prime\prime}\text{ are themost efficient Nash equilibria of }\mathcal{G}^{\prime},W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})\right]$ (59)
@@
$\sum_{j=1}^{N}\left(\gamma_{j,\pi^{\prime}_{j}}-\gamma_{j,\pi^{\prime\prime}_{j}}\right)=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi^{\prime}_{j}}w_{j,\pi^{\prime}_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi^{\prime}_{j^{\prime}}=\pi^{\prime}_{j}]w_{j^{\prime},\pi^{\prime}_{j}}}-\frac{\hat{\mu}_{j,\pi^{\prime\prime}_{j}}w_{j,\pi^{\prime\prime}_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi^{\prime\prime}_{j^{\prime}}=\pi^{\prime\prime}_{j}]w_{j^{\prime},\pi^{\prime\prime}_{j}}}\right).$ (60)
Note that this occurs with probability 00 because $\gamma_{j,k}$ are sampled uniformly from $[0,\Gamma]$ and the right-hand side of the above equation is a constant when $\bm{\pi}^{\prime}$ and $\bm{\pi}^{\prime\prime}$ are fixed. As a result, we have $\forall\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N}$, $\mathbb{P}[W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})]=0$ and the left-hand side of Equation (59) is 00. Now the claim follows. ∎
C.4 Proof of Theorem 4.2
@@
Let $\mathcal{Z}$ be the space of all possible state of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$. In addition, define $\mathcal{D}$ as the set of all state in the Markov process where all players are discontent.
$\mathcal{D}=\left\{Z=\{(m_{j},a_{j},u_{j})\}_{j=1}^{N}\in\mathcal{Z}:\forall j\in[N],m_{j}=D\right\}.$ (61)
We first note that similar to [Pradelski and Young, 2012], the transition probability from any state $Z_{1}\in\mathcal{Z}\backslash\mathcal{D}$ to $Z_{2}\in\mathcal{D}$ and the transition probability from any state $Z_{2}\in\mathcal{D}$ to $Z_{1}\in\mathcal{Z}\backslash\mathcal{D}$ are independent of the $Z_{2}$. As a result, we can view all states in $\mathcal{D}$ as a single state. The transition probability from $\mathcal{D}$ to other state $Z\in\mathcal{Z}\backslash\mathcal{D}$ is the same as the probability from any state $Z_{1}\in\mathcal{D}$ to $Z$. We slightly abuse the notation here to denote the Markov process as $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ after combining all the states in $\mathcal{D}$ as a single state.
@@
$\lim_{\epsilon\rightarrow 0^{+}}\mathcal{M}_{\epsilon}(Z\rightarrow Z^{\prime})=\mathcal{M}_{0}(Z\rightarrow Z^{\prime})$ (62)
@@
$0<\lim_{\epsilon\rightarrow 0^{+}}\frac{\mathcal{M}_{\epsilon}(Z\rightarrow Z^{\prime})}{\epsilon^{r(Z\rightarrow Z^{\prime})}}<\infty$ (63)
@@
$\lim_{\epsilon\rightarrow 0^{+}}p_{Z}(\epsilon)>0.$ (64)
@@
$Z^{*}=\{(m_{j}^{*},a_{j}^{*},u_{j}^{*})\}_{j\in[N]}\text{ with }\forall j\in[N],m_{j}^{*}=C,a_{j}^{*}=\pi_{j}^{*},u_{j}^{*}=\mathcal{U}^{\prime}_{j}(\bm{\pi}^{*})$ (65)
@@
(b) We further demonstrate the functions $F(\cdot)$ and $G(\cdot)$ in Equation (9) meet certain conditions.
For any $j\in[N]$, define the set $V_{j}=\left\{\mathcal{U}^{\prime}_{j}(\bm{\pi}):\bm{\pi}\in[K]^{N}\right\}$ where $\mathcal{U}^{\prime}_{j}(\bm{\pi})$ is provided in Equation (7). As a result,
$\forall u^{\prime}\in V_{j},u^{\prime}\leq\max_{k}(\hat{\mu}_{j,k}+\gamma_{j,k})\leq 1+\Gamma,\quad\text{and}\quad\forall u^{\prime},u^{\prime\prime}\in V_{j},\left|u^{\prime}-u^{\prime\prime}\right|\leq\max_{u^{\prime}\in V_{j}}u^{\prime}\leq 1+\Gamma.$ (66)
Therefore,
$\forall u^{\prime}\in V_{j},\quad F(u^{\prime})=-\frac{u^{\prime}}{(4+3\Gamma)N}+\frac{1}{3N}>-\frac{1+\Gamma}{(4+3\Gamma)N}+\frac{1}{3N}>0\text{ and }F(u^{\prime})=-\frac{u^{\prime}}{(4+3\Gamma)N}+\frac{1}{3N}<\frac{1}{3N}<\frac{1}{2N}.$ (67)
In addition,
$\forall u^{\prime},u^{\prime\prime}\in V_{j},u^{\prime}\geq u^{\prime\prime},\quad G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}\geq-\frac{1+\Gamma}{4+3\Gamma}+\frac{1}{3}>0\text{ and }G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}<\frac{1}{3}<\frac{1}{2}.$ (68)
(c) Finally, we could prove this lemma according to Theorem 1 in [Pradelski and Young, 2012]. Note that our Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ follow exactly the one proposed in [Pradelski and Young, 2012] and the only modifications are to ensure that the payoff calculation for each player follows the game $\mathcal{G}^{\prime}$. By part (a) and (b) of our proof, the conditions of Theorem 1 in [Pradelski and Young, 2012] hold. As a result, since $\mathcal{G}^{\prime}$ have at least one PNE and the most efficient PNE is unique with probability $1$ according to Theorem 4.1, the only stochastically stable state is $Z^{*}$ in Equation (65). As a result, when $\epsilon\rightarrow 0$, $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})\rightarrow 1$. Now the claim follows. ∎
@@
Suppose the conditions in Theorem 4.2 hold. Let $w_{s}$ be the set of rounds in the learning phase with length $c_{2}s^{\eta}$ and $Z^{*}$ be the unique stochastically stable state of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ given by Equation (65). Let $Z(t)$ be status of the Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ at round $t$. Define the counter $H_{s}$ and the event $E_{2s}$ as
[…]
@@
$\mathbb{P}[E_{2s}]\geq 1-\left(\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\right)^{s},$ (70)
@@
$\mathbb{P}\left[H_{s}\leq(1-\alpha)p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}\right]\leq\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right),$ (71)
Note that $\{H_{s}\}$ are independent. As a result, according to the Chernoff bound, we have
$\forall\xi>0,\quad\mathbb{P}\left[\bar{E}_{2s}\right]=\mathbb{P}\left[\sum_{i=1}^{s}H_{i}\leq\frac{L_{s}}{2}\right]\leq\exp\left(\frac{\xi L_{s}}{2}\right)\mathbb{E}\left[\prod_{i=1}^{s}\exp\left(-\xi H_{i}\right)\right]=\exp\left(\frac{\xi L_{s}}{2}\right)\prod_{i=1}^{s}\mathbb{E}\left[\exp\left(-\xi H_{i}\right)\right].$ (72)
For every $i$ such that $1\leq i\leq s$, by Equation (71) we have
$\,\mathbb{E}[\exp(-\xi H_{i})]$ (73)
@@
$\mathbb{E}[\exp(-\xi H_{i})]\leq\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}i^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right).$ (74)
@@
$\mathbb{P}[\bar{E}_{2s}]$ $\leq\exp\left(\frac{\alpha^{2}L_{s}}{144(1-\alpha)A(\epsilon;\mathcal{G}^{\prime})}\right)\prod_{i=1}^{s}\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}i^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right)$ (75)
@@
$L_{s}=\sum_{i=1}^{s}c_{2}i^{\eta}\geq c_{2}\int_{0}^{s-1}i^{\eta}\mathrm{d}i=\frac{c_{2}(s-1)^{\eta+1}}{\eta+1}\geq c_{2}\left(\frac{s}{2}\right)^{\eta+1}$ (76)
@@
$\mathbb{P}[\bar{E}_{2s}]\leq\left(\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\right)^{s}.$ (77)
@@
$E_{1}=\left\{\forall j\in[N],k\in[K],\left|\hat{\mu}_{j,k}-\mu_{k}\right|<\frac{\delta}{4K}\right\}.$ (78)
@@
$\mathbb{P}\left[\left|\hat{\mu}_{j,k}-\mu\right|>\frac{\delta}{4K}\right]\leq 2\exp\left(-\frac{c_{1}\delta^{2}}{8K^{2}}\right)\leq\frac{2}{T}.$ (79)
@@
$\mathbb{P}(E_{1})\geq 1-\sum_{j=1}^{N}\sum_{k=1}^{K}\mathbb{P}\left[\left|\hat{\mu}_{j,k}-\mu\right|>\frac{\delta}{4K}\right]\geq 1-\frac{2NK}{T}.$ (80)
@@
(b) Consider the learning and exploitation phase. Under the clean event $E_{1}$, according to Theorem 4.1, we have that the most efficient Nash equilibrium $\bm{\pi}^{*}$ of game $\mathcal{G}^{\prime}$ is also a most efficient equilibrium of game $\mathcal{G}$. The solution is also unique with probability $1$. As a result, according to Lemma C.5, for any $10<\alpha<$ 1 there exists a small enough $\epsilon$ such that the state $Z^{*}$ in Equation (65) is the unique stationary distribution of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ and $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$.
@@
$\,\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]$ (81)
@@
$\forall s>S_{0},\quad B(\epsilon;\mathcal{G}^{\prime})\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\leq\frac{1}{2}.$ (82)
@@
$\,\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]$ (83)
@@
$\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]\leq 2\mu_{\max}\left(\frac{c_{2}(S+1)^{\eta+1}}{\eta+1}+c_{3}S\right)\leq O\left(\mu_{\max}\left(c_{2}\log^{1+\eta}T+c_{3}\log T\right)\right).$ (84)
@@
$\mathbb{E}[\operatorname{Reg}_{j}(T)]$ $=\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{init}}(T)\right]+\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\right]$ (85)
@@
$\mathbb{E}[\operatorname{NonPNE}(T)]$ $=\mathbb{E}\left[\operatorname{NonPNE}^{\text{init}}(T)\right]+\mathbb{E}\left[\operatorname{NonPNE}^{\text{learn}}(T)\right]$ (86)

### v2

@@ Appendix C Omitted Proofs
$W_{\mathcal{G}^{\prime}}(\bm{\pi})=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi_{j}}w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}+\gamma_{j,\pi_{j}}\right)\geq\sum_{j=1}^{N}\frac{(\mu_{\pi_{j}}-\Delta)w_{j,\pi_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=\pi_{j}]w_{j^{\prime},\pi_{j}}}=\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}-k\Delta\geq\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}-K\Delta.$
@@
$W_{\mathcal{G}^{\prime}}(\tilde{\bm{\pi}})-W_{\mathcal{G}^{\prime}}(\bm{\pi})\geq\sum_{k^{\prime}=1}^{\tilde{k}}\mu_{k^{\prime}}-K\Delta-\left(\sum_{k^{\prime}=1}^{k}\mu_{k^{\prime}}+K\Delta+N\Gamma\right)\geq\mu_{\tilde{k}}-2K\Delta-N\Gamma.$
@@
$W_{\mathcal{G}^{\prime}}(\tilde{\bm{\pi}})-W_{\mathcal{G}^{\prime}}(\bm{\pi})\geq\mu_{\tilde{k}}-2K\Delta-N\Gamma\geq\delta-2K\cdot\frac{\delta}{4K}-N\cdot\frac{\delta}{4N}=\frac{\delta}{4}>0.$
@@
$\,\mathbb{P}\left[\exists\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N},\bm{\pi}^{\prime}\text{ and }\bm{\pi}^{\prime\prime}\text{ are themost efficient Nash equilibria of }\mathcal{G}^{\prime},W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})\right]$
@@
$\sum_{j=1}^{N}\left(\gamma_{j,\pi^{\prime}_{j}}-\gamma_{j,\pi^{\prime\prime}_{j}}\right)=\sum_{j=1}^{N}\left(\frac{\hat{\mu}_{j,\pi^{\prime}_{j}}w_{j,\pi^{\prime}_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi^{\prime}_{j^{\prime}}=\pi^{\prime}_{j}]w_{j^{\prime},\pi^{\prime}_{j}}}-\frac{\hat{\mu}_{j,\pi^{\prime\prime}_{j}}w_{j,\pi^{\prime\prime}_{j}}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi^{\prime\prime}_{j^{\prime}}=\pi^{\prime\prime}_{j}]w_{j^{\prime},\pi^{\prime\prime}_{j}}}\right).$
Note that this occurs with probability 00 because $\gamma_{j,k}$ are sampled uniformly from $[0,\Gamma]$ and the right-hand side of the above equation is a constant when $\bm{\pi}^{\prime}$ and $\bm{\pi}^{\prime\prime}$ are fixed. As a result, we have $\forall\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N}$, $\mathbb{P}[W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})]=0$ and the left-hand side of Section C.4 is 00. Now the claim follows. ∎
C.5 Proof of Theorem 4.2
@@
Let $\mathcal{Z}$ be the space of all possible states of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$. In addition, define $\mathcal{D}$ as the set of all states in the Markov process where all players are discontent.
$\mathcal{D}=\left\{Z=\{(m_{j},a_{j},u_{j})\}_{j=1}^{N}\in\mathcal{Z}:\forall j\in[N],m_{j}=D\right\}.$
We first note that similar to [Pradelski and Young, 2012], the transition probability from any state $Z_{1}\in\mathcal{Z}\backslash\mathcal{D}$ to $Z_{2}\in\mathcal{D}$ and the transition probability from any state $Z_{2}\in\mathcal{D}$ to $Z_{1}\in\mathcal{Z}\backslash\mathcal{D}$ are independent of the $Z_{2}$. As a result, we can view all states in $\mathcal{D}$ as a single state. The transition probability from $\mathcal{D}$ to another state $Z\in\mathcal{Z}\backslash\mathcal{D}$ is the same as the probability from any state $Z_{1}\in\mathcal{D}$ to $Z$. We slightly abuse the notation here to denote the Markov process as $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ after combining all the states in $\mathcal{D}$ as a single state.
@@
$\lim_{\epsilon\rightarrow 0^{+}}\mathcal{M}_{\epsilon}(Z\rightarrow Z^{\prime})=\mathcal{M}_{0}(Z\rightarrow Z^{\prime})$
@@
$0<\lim_{\epsilon\rightarrow 0^{+}}\frac{\mathcal{M}_{\epsilon}(Z\rightarrow Z^{\prime})}{\epsilon^{r(Z\rightarrow Z^{\prime})}}<\infty$
@@
$\lim_{\epsilon\rightarrow 0^{+}}p_{Z}(\epsilon)>0.$
@@
$Z^{*}=\{(m_{j}^{*},a_{j}^{*},u_{j}^{*})\}_{j\in[N]}\text{ with }\forall j\in[N],m_{j}^{*}=C,a_{j}^{*}=\pi_{j}^{*},u_{j}^{*}=\mathcal{U}^{\prime}_{j}(\bm{\pi}^{*})$
@@
(b) We further demonstrate the functions $F(\cdot)$ and $G(\cdot)$ in Eq. 7 meet certain conditions.
For any $j\in[N]$, define the set $V_{j}=\left\{\mathcal{U}^{\prime}_{j}(\bm{\pi}):\bm{\pi}\in[K]^{N}\right\}$ where $\mathcal{U}^{\prime}_{j}(\bm{\pi})$ is provided in Eq. 6. As a result,
$\forall u^{\prime}\in V_{j},u^{\prime}\leq\max_{k}(\hat{\mu}_{j,k}+\gamma_{j,k})\leq 1+\Gamma,\quad\text{and}\quad\forall u^{\prime},u^{\prime\prime}\in V_{j},\left|u^{\prime}-u^{\prime\prime}\right|\leq\max_{u^{\prime}\in V_{j}}u^{\prime}\leq 1+\Gamma.$
Therefore, $\forall u^{\prime}\in V_{j},\quad$,
$F(u^{\prime})=-\frac{u^{\prime}}{(4+3\Gamma)N}+\frac{1}{3N}>-\frac{1+\Gamma}{(4+3\Gamma)N}+\frac{1}{3N}>0\text{ and }F(u^{\prime})=-\frac{u^{\prime}}{(4+3\Gamma)N}+\frac{1}{3N}<\frac{1}{3N}<\frac{1}{2N}.$
In addition, $\forall u^{\prime},u^{\prime\prime}\in V_{j},u^{\prime}\geq u^{\prime\prime}$,
$G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}\geq-\frac{1+\Gamma}{4+3\Gamma}+\frac{1}{3}>0\text{ and }G(u^{\prime}-u^{\prime\prime})=-\frac{u^{\prime}-u^{\prime\prime}}{4+3\Gamma}+\frac{1}{3}<\frac{1}{3}<\frac{1}{2}.$
(c) Finally, we could prove this lemma according to Theorem 1 in [Pradelski and Young, 2012]. Note that our Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ follow exactly the one proposed in [Pradelski and Young, 2012] and the only modifications are to ensure that the payoff calculation for each player follows the game $\mathcal{G}^{\prime}$. By part (a) and (b) of our proof, the conditions of Theorem 1 in [Pradelski and Young, 2012] hold. As a result, since $\mathcal{G}^{\prime}$ have at least one PNE and the most efficient PNE is unique with probability $1$ according to Theorem 4.1, the only stochastically stable state is $Z^{*}$ in Lemma C.5. As a result, when $\epsilon\rightarrow 0$, $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})\rightarrow 1$. Now the claim follows. ∎
@@
Suppose the conditions in Theorem 4.2 hold. Let $w_{s}$ be the set of rounds in the learning phase with length $c_{2}s^{\eta}$ and $Z^{*}$ be the unique stochastically stable state of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ given by Lemma C.5. Let $Z(t)$ be status of the Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ at round $t$. Define the counter $H_{s}$ and the event $E_{2s}$ as
$H_{s}=\sum_{t\in w_{s}}\mathbb{I}[Z(t)=Z^{*}]\quad\text{and}\quad E_{2s}=\left\{\sum_{i=1}^{s}H_{i}>\frac{L_{s}}{2}\right\},L_{s}=\sum_{i=1}^{s}c_{2}i^{\eta}.$
@@
$\mathbb{P}[E_{2s}]\geq 1-\left(\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\right)^{s},$
@@
$\mathbb{P}\left[H_{s}\leq(1-\alpha)p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}\right]\leq\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}s^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right),$
Note that $\{H_{s}\}$ are independent. As a result, according to the Chernoff bound, we have $\forall\xi>0$,
$\mathbb{P}\left[\bar{E}_{2s}\right]=\mathbb{P}\left[\sum_{i=1}^{s}H_{i}\leq\frac{L_{s}}{2}\right]\leq\exp\left(\frac{\xi L_{s}}{2}\right)\mathbb{E}\left[\prod_{i=1}^{s}\exp\left(-\xi H_{i}\right)\right]=\exp\left(\frac{\xi L_{s}}{2}\right)\prod_{i=1}^{s}\mathbb{E}\left[\exp\left(-\xi H_{i}\right)\right].$
For every $i$ such that $1\leq i\leq s$, by Section C.5 we have
$\,\mathbb{E}[\exp(-\xi H_{i})]$
@@
$\mathbb{E}[\exp(-\xi H_{i})]\leq\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}i^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right).$
@@
$\mathbb{P}[\bar{E}_{2s}]$ $\leq\exp\left(\frac{\alpha^{2}L_{s}}{144(1-\alpha)A(\epsilon;\mathcal{G}^{\prime})}\right)\prod_{i=1}^{s}\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})c_{2}i^{\eta}}{72A(\epsilon;\mathcal{G}^{\prime})}\right)$
@@
$L_{s}=\sum_{i=1}^{s}c_{2}i^{\eta}\geq c_{2}\int_{0}^{s-1}i^{\eta}\mathrm{d}i=\frac{c_{2}(s-1)^{\eta+1}}{\eta+1}\geq c_{2}\left(\frac{s}{2}\right)^{\eta+1}$
@@
$\mathbb{P}[\bar{E}_{2s}]\leq\left(\left(1+\frac{C(\epsilon;\mathcal{G}^{\prime})}{\sqrt{p_{\mathcal{D}}(\epsilon;\mathcal{G}^{\prime})}}\right)\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\right)^{s}.$
@@
$E_{1}=\left\{\forall j\in[N],k\in[K],\left|\hat{\mu}_{j,k}-\mu_{k}\right|<\frac{\delta}{4K}\right\}.$
@@
$\mathbb{P}\left[\left|\hat{\mu}_{j,k}-\mu\right|>\frac{\delta}{4K}\right]\leq 2\exp\left(-\frac{c_{1}\delta^{2}}{8K^{2}}\right)\leq\frac{2}{T}.$
@@
$\mathbb{P}(E_{1})\geq 1-\sum_{j=1}^{N}\sum_{k=1}^{K}\mathbb{P}\left[\left|\hat{\mu}_{j,k}-\mu\right|>\frac{\delta}{4K}\right]\geq 1-\frac{2NK}{T}.$
@@
(b) Consider the learning and exploitation phase. Under the clean event $E_{1}$, according to Theorem 4.1, we have that the most efficient Nash equilibrium $\bm{\pi}^{*}$ of game $\mathcal{G}^{\prime}$ is also a most efficient equilibrium of game $\mathcal{G}$. The solution is also unique with probability $1$. As a result, according to Lemma C.5, for any $10<\alpha<$ 1 there exists a small enough $\epsilon$ such that the state $Z^{*}$ in Lemma C.5 is the unique stationary distribution of $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ and $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})>1/(2(1-\alpha))$.
@@
$\,\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]$
@@
$\forall s>S_{0},\quad B(\epsilon;\mathcal{G}^{\prime})\exp\left(-\frac{\alpha^{2}c_{2}}{144A(\epsilon;\mathcal{G}^{\prime})}\left(p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})-\frac{1}{2(1-\alpha)}\right)\left(\frac{s}{2}\right)^{\eta}\right)\leq\frac{1}{2}.$
@@
$\,\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]$
@@
$\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\mid E_{1}\right]\leq 2\mu_{\max}\left(\frac{c_{2}(S+1)^{\eta+1}}{\eta+1}+c_{3}S\right)\leq O\left(\mu_{\max}\left(c_{2}\log^{1+\eta}T+c_{3}\log T\right)\right).$
@@
$\mathbb{E}[\operatorname{Reg}_{j}(T)]$ $=\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{init}}(T)\right]+\mathbb{E}\left[\operatorname{Reg}_{j}^{\text{learn}}(T)\right]$
@@
$\mathbb{E}[\operatorname{NonPNE}(T)]$ $=\mathbb{E}\left[\operatorname{NonPNE}^{\text{init}}(T)\right]+\mathbb{E}\left[\operatorname{NonPNE}^{\text{learn}}(T)\right]$